OpenStudy (anonymous):

if sin theta=8 over 11, what is sec theta

OpenStudy (anonymous):



OpenStudy (anonymous):

there is an angle \(\theta \) with \(\sin(\theta )=\frac{8}{11}\) all you need is \(a\) which you find via pythgoras

OpenStudy (anonymous):

then you can compute any ratio you like, including \(\sec(\theta)=\frac{a}{8}\)

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

yw